OpenStudy (anonymous):

Add and simplify: x + 1/x^2 + 3/4x

OpenStudy (anonymous):

@phi ?

OpenStudy (phi):

can you use the equation editor, or at least use parens for this one ?

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

\[\frac{ x+1 }{ x^2 } + \frac{ 3 }{4x}\]

OpenStudy (phi):

can you multiply the first fraction by 4/4 ? and the 2nd fraction by x/x ?

OpenStudy (anonymous):

ya uh... \[\frac{ 4x+1 }{ 4x^2 } + \frac{ 3x }{ 4x^2 }\]

OpenStudy (anonymous):

or is it +4

OpenStudy (phi):

close. the first fraction, up top is (x+1)*4 you can then "distribute" the 4: 4 times the x and 4 times the 1 you get 4x+4 (not 4x+1) after you fix that, you can add the tops, because the fractions have the same bottoms

OpenStudy (anonymous):

ok then you get\[\frac{ 7x +4 }{ 8x^2 }\] right?

OpenStudy (phi):

almost. when adding fractions with the same bottom... the sum has *the same* bottom. You don't do anything to the bottom... it stays 4x^2

OpenStudy (anonymous):

ok yeah ok so it is 7x+4/4x^2?

OpenStudy (phi):

if we had numbers 1/3 + 1/3 we would add the tops to get 1+1=2 and keep the same bottom 1/3 + 1/3 = 2/3 we use the same rule even if the bottom has letters

OpenStudy (anonymous):

ok that makes sense. Thanks!